OpenStudy (anonymous):

2 Batches of soup serves 8 people. Lonie needsx to make enought soup for 120 people. if you need 2 tomatoes for Each batch hoe many tomatoes will she need?

OpenStudy (saifoo.khan):

Ratio Method.

OpenStudy (anonymous):

Would that be 120 + 120 = 240?

OpenStudy (anonymous):

What is ratio method?

OpenStudy (diyadiya):

1 batch = 4 ppl 30batch=120 ppl 2tomats for 1 batch 60 tomts for 30 batch

OpenStudy (anonymous):

Is the answer 240?

OpenStudy (anonymous):

it's 60

OpenStudy (diyadiya):

no its 60

OpenStudy (saifoo.khan):

Yea!!

OpenStudy (anonymous):

how did you get 60?

OpenStudy (saifoo.khan):

THere are 30 batches.. okay? now: tomatos : Batch 2 : 1 x : 30 Cross multiply.

OpenStudy (anonymous):

(: Thanks!

OpenStudy (saifoo.khan):

(: Welks!

OpenStudy (anonymous):

did you understand why there are 30 batches ? :)

OpenStudy (anonymous):

Yups! (((:

OpenStudy (diyadiya):

2 batch = 8ppl 1 batch = 8/2=4ppl xbatch*4=120ppl =120/4=30batch